OpenStudy (anonymous):

Austin takes 1 minute and 45 seconds to run three-quarters of a circular track. His rate of motion is π/ radians per second.

OpenStudy (uri):

distance he traveled in radian=2*pi*3/4. @alexrico

OpenStudy (anonymous):

\[\frac{3\ 2 \pi }{4\ 105}=\frac{1}{70} \pi \text{ radians} \text{ per second} \]